# Evaluating the Scalability of Binary and Ternary CNN Workloads on RRAM-based Compute-in-Memory Accelerators

### v1

@@ III-B ADC Calibration
III-C Energy Model
Previous work has shown that the energy consumed for MVM operation on RRAM crossbars is strongly dependent on the operand encoding, matrix mapping and tiling schemes [5, 2].
We implemented an additive, statistical energy model that relies on the technology-dependent reference values provided by Accelergy [17]. The model is additive, as operations in the workload are counted, multiplied by reference values from the technology library, and added together. Workload mapping and operand distribution statistics are considered via scaling factors.
Our model considers only the crossbar array and the mixed-signal peripheral circuits, specifically the Row Drivers (RD) and the ADC. Reference energy values for RD $E_{\text{RD}}$ are drawn from NeuroSim [14], while the ADC energy ($E_{\text{ADC}}$) comes from the analytical model in [1]. RRAM programming costs are not considered, as we are only interested in the run-time energy costs.
As the workload is partitioned into tiles mapped to crossbars, we first derive the energy for each tile as:
@@ III-C Energy Model
where $O_{t}$ is the number of MVM operations done for tile $t$, $N_{t}$ is the number of active rows, $M_{t}$ is the number of active columns, $\overline{\mathbf{x}}_{t}\in[0,1]$ is the average input value, and $\overline{\mathbf{g}}_{t}$ is the average cell conductance. The last term represents the energy spent in the memristive cells, which depends on the read voltage $V_{\text{R}}$ and effective pulse length $T_{\text{R}}$ [5].
@@ IV-A Exploration Framework
To validate our work, we built an exploration framework consisting of two main components: a TVM-based compiler and a CIM-core simulator. The TVM-based compiler translates convolutional and dense layers to tiled MVM operations and offloads these to the CIM-core simulator.
The CIM-core simulator integrates all models presented in this work: the parasitic-aware current solver from Section III-A, the ADC model with adjustable scale from Section III-B and the energy model from Section III-C. It supports all operand mappings described in II-B and features workload statistics collection utilities. All simulation components are implemented in Python and rely heavily on the Numpy[8] library to accelerate matrix operations.

### v2

@@ III-B ADC Calibration
III-C CIM-Core Energy Model
Previous work has shown that the energy consumed by MVM operations on RRAM crossbars is strongly dependent on the operand encoding, matrix mapping and tiling schemes [5, 2].
We implemented an additive, statistical energy model that uses the crossbar energy model from [5]. Additionally, we consider the main mixed-signal peripheral circuits in the CIM-Core depicted in Fig. 1a: Row Drivers (RD) and ADC. Reference energy values for RD $E_{\text{RD}}$ are drawn from NeuroSim [15], while for the ADC ($E_{\text{ADC}}$) these come from the analytical tool in [1]. RRAM programming costs are not considered, as we are only interested in the run-time energy costs.
As explained in Section II-B, convolution and dense layers in CNN workloads are partitioned into tiles and mapped to crossbars. We estimate the energy for each tile as:
@@ III-C CIM-Core Energy Model
where $O_{t}$ is the number of MVM operations done for tile $t$, $N_{t}$ is the number of active rows, $M_{t}$ is the number of active columns, $\overline{\mathbf{x}}_{t}\in[0,1]$ is the average input value, and $\overline{\mathbf{g}}_{t}$ is the average cell conductance. The last term in 8 represents the energy spent in the memristive cells, which depends on the read voltage $V_{\text{R}}$ and effective pulse length $T_{\text{R}}$ [5]. The scaling factors $\overline{\mathbf{x}}_{t}$ and $\overline{\mathbf{g}}_{t}$ account for operand distribution statistics and the selected encoding scheme.
@@ IV-A Exploration Framework
To validate our work, we created an extended version of CIM-Explorer [14], an exploration framework consisting a TVM-based compiler and a CIM-core simulator. The TVM-based compiler translates convolutional and dense layers to tiled MVM operations and offloads these to the CIM-core simulator [14].
The extended CIM-core simulator integrates all models presented in this work: the parasitic-aware current solver from Section III-A, the ADC model with adjustable scale from Section III-B and the energy model from Section III-C. It supports all operand mappings described in II-B and features workload statistics collection utilities. All simulation components are implemented in Python and rely heavily on the Numpy[8] library to accelerate matrix operations.